September 13, 2013 Installation
Meeting Time and Place
Friday, September 13, 2013 at 1:30 PM At Jefferson Lab, the meeting will be held in F326
To connect from the outside, please use ESNET
- ) ESNET: 8542553
- )To connect by telephone (supported on request only) dial:
- You can look up toll-free number at http://www.readytalk.com/intl
- US and Canada: (866)740-1260 (toll free)
- International: (303)248-0285 (toll call)
- enter access code followed by the # sign: 3421244#
You can view the computer desktop in the meeting room at JLab via the web.
- Go to http://esnet.readytalk.com
- In the "join a meeting" box enter the Hall D code: 3421244
- Fill in the participant registration form.
Action Items From Last Meeting
- Magnet Status
- Impact of Magnet work on installation.
- Updates to the Mechanical Installation Schedule ( Sept Rendition of Install Schedule) (Tim)
- Electronics Installation Schedule (Fernando) [| Schedule]
- Installation Updates (Not a construction report)
- Counting House Infrastructure
- Trigger and DAQ Installation
- Slow Control Installation
- Tagger Magnet Installation
- BCAL Installation Downstream View, Upstream View, Closeup Upstream View
- FCAL Installation
- FDC Installation
- CDC Installation
- TOF Installation
- Start Counter Installation
- Microscope Installation
- Tagger Array Installation
- Pair Spectrometer Installation
Present: JLab, IU, CMU, Regina, FSU
Rebaselining is approved, but there is more scheduled manpower than is available in the coming two months.
Directors review of the magnet at the end of September. There is a difference in opinions in what caused the quench. An independent internal report (by Lee Harwood) concluded that there was a short in the magnet. If true, this suggests that the magnet is much more difficult to use. There is some debate about whether the observed behavior could be due to a short. Many think it is refrigeration problem alone, which is solvable with an upgraded refrigerator. Discussion is ongoing and a final report needs to be prepared and presented.
Review of schedule:
Tim posted a revised installation schedule. Next critical path item is to reassemble the upstream platform for the FDC installation.
Tagger microscope is missing from this schedule.
Sasha expects delivery of beamline instruments on time. There will be some delay with the tagger hodoscope.
Current plans are to have beam in tagger hall only in April. We will want to have beam instrumentation available.
We should plan to have a commissioning target available in April. Looking ahead there was a projection that the scheduled Start Counter installation may be tight.Our goal should be to have start counter and commissioning target ready in April and fallback can be to insert commissioning target w/o start counter if need be.
Electronics schedule (Fernando):
Tagger cabling should be advanced to make things ready for April.
Large lag between installation of start counter and cabling of start counter.
System checkout for FCAL can be advanced.
Current resource loaded schedule assumes only lab technical staff. Items can be advanced in the schedule with the addition of contributed labor from the universities.
Counting House and Networking (Elliot): Infrastructure is in good shape. Tagger hall networking will be done next week. Interim network is being installed in the hall.
Trigger and DAQ (Sascha): Two trigger crates are checked and ready to go to hall. Need optical fiber bundles (to be ordered in October). Installation on schedule. DAQ plan is to have CODA3 ready for any DAQ crate.
Slow Control (Elliot): Installation of hardware in the hall is beginning for control and monitoring. Start with networking in the hall near the FCAL for control system.
Tagger Magnet (Tim): Working to power it up and map it.
BCAL: it's in!
FCAL: 3/4 of detector cabled and getting ready for darkroom instrumentation.
FDC (Tim): Some discussion is needed to understand how to align the FDC to the BCAL. Plan to survey the BCAL Monday to understand the position and set the rails to center of BCAL. Then further discussion on alignment will happen once data is available.
CDC (Curtis): Plan to deliver late this year. There may be room in the EEL building to work with the detector.
TOF (Paul): Information on installation details is linked to the agenda. Password and username is same as that used for document database. Tim will be sure space is available.
Start Counter (Mark): If we want the Start Counter in April then we need to be sure electronics is ready.
Microscope, tagger array, pair spectrometer (Eugene, Sascha): Some delay due to construction delay and anticipate February installation. Dan Physik will come inspect power supply for pair spectrometer next week. Damage may be mechanical only.